find the y-intercept of a line with the equation
y= -3/13x + 13/6

Answer by nerdybill(7384)

find the y-intercept of a line with the equation
y= -3/13x + 13/6
.
"slope-intercept" form of a line is:
y = mx + b
where
m is the slope
b is the y-intercept at (0,b)
.
Your equation:
y= -3/13x + 13/6
is already in "slope-intercept" form
so, from inspection then
the y-intercept is at (0, 13/6)